OpenStudy (vespersaviour):

What is the equation of the line? A. y=-3/2x B.y=-2/3x C.y=2/3x D.y=3/2x

OpenStudy (vespersaviour):

@knov

563blackghost (563blackghost):

So to find the slope we would pick two points on the graph...lets use (0,0) and (3,2)...we will use the slope formula to find the slope.... \(\Huge{slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}\)

OpenStudy (vespersaviour):

\[\frac{ -2 }{ -3}\]??

563blackghost (563blackghost):

Not quite. We would set up as.. \(\Huge{slope=\frac{2-0}{3-0}}\)

OpenStudy (vespersaviour):

so it would be \[\frac{ 2 }{ 3 }\] ?

563blackghost (563blackghost):

Yup ^^

OpenStudy (vespersaviour):

Thank you!

563blackghost (563blackghost):

np ^^